# Rewarding Curse: Analyze and Mitigate Reward Modeling Issues for LLM Reasoning

### v1

@@ 8 Conclusion
In this work, we focus on analyzing key factors that influence the reward model’s performance in reasoning tasks. We find that low question difficulty, large sampling number, and high search diversity can lead to issues in RM-based inference, with in-depth explanations provided. After analyzing these issues, we design a new inference algorithm called OCTS to mitigate them. Experimental results demonstrate that our method is effective in enhancing RM-based reasoning capabilities.
Limitations
Although our work conducts an in-depth analysis and mitigating in RM-based reasoning, it has several limitations. Firstly, we refrain from analyzing the problem of reward models in more reasoning tasks such as commonsense and logic because the LLM’s performance on these tasks is already sufficiently strong. Secondly, in this work, we focus solely on investigating the issues present during the inference phase, while potential problems that may arise during the training of RM are not explored. This is because the training process involves significant randomness and a large number of parameters, making it challenging to pinpoint the key factors that affect RM’s performance. We leave these limitations as our future work to explore.
Appendix A Performance of Selected RMs
To demonstrate that the RM issues identified in our experiments in Section §2.1 are not due to the selected RM’s inherently low discriminative abilities, here we present the performance of our RM. For the two ORMs (e.g. ArmoRM-Llama3-8B and Skywork-Reward-Llama-3.1-8B), we report their performance on RewardBench (Lambert et al., 2024) compared to other baselines in Table 2. For the two PRMs (e.g. Math-Shepherd-Mistral-7B-PRM and Skywork-o1-Open-PRM-Qwen-2.5-7B), we report their performance on ProcessBench (Lambert et al., 2024) compared to other baselines in Table 3. From them, we can get that the performance of these models on relevant benchmarks is comparable to the advanced LLMs (e.g. gpt4), hence they are representative.
Appendix B Additional Overall Experiments
@@
In addition, we also use the consistency of the final answer output by the policy model itself as the source of the reward, denoted as ‘Self’. The results are demonstrated in Figure 12. We can conclude that: (1) Even with the MCTS framework, the improvement in model reasoning brought by the RM is still minimal, further validating our conclusions in §2.1. (2) In Skywork and Skyworko1, the average performance of Reward is the best among all scoring functions. Therefore, in the MCTS-related experiments presented in the main text, we default to using it as the scoring function $f$.
Secondly, we focus on math reasoning in the main text, here we repeat our experiments on other types of reasoning tasks. Specifically, for math reasoning, we select another dataset: AQuA (Ling et al., 2017). For commonsense reasoning, we select WinoGrande (WINO) (Sakaguchi et al., 2020) and CSQA (Talmor et al., 2019); For logical reasoning, we select ProofWriter (Tafjord et al., 2021) and ProntoQA (Saparov and He, 2023) The results are demonstrated in Figure 13, 14, 15, 16 and 17. Lastly, we only use discriminative RM in the main text. All of these results are consistent with the conclusion in the main text.
Appendix C Additional Experiments on Question Difficulty Approximation
@@
We classify the problems according to the difficulty levels as outlined in the main text and calculate the above three metrics across different levels of problem difficulty to compare the degree of correlation. The results are illustrated in Figure 18, 19 and 20. We can observe that, comparatively, the Count function is most directly proportional to difficulty. Therefore, we use this function to estimate difficulty when designing the OCTS method in §5.1.
Appendix D Additional Experiments across Different Difficulty Levels
In the main text, we only analyze the impact of question difficulty on the MATH dataset. To demonstrate the generalizability of our conclusions, we repeat this experiment on GSM8k (Cobbe et al., 2021) and Olympiadbench (He et al., 2024). The former dataset contains 8.5K linguistically diverse elementary school math problems designed to evaluate arithmetic reasoning consistency, while the latter is an Olympiad-level bilingual multimodal scientific benchmark. Compared to MATH, the former is simpler, while the latter is more challenging. The results are illustrated in Table 4, 5 and 6. We can observe that the issues identified in Cl.1 are prevalent across various reasoning datasets.
Appendix E Diversity Experiment on Exploration Constant
@@
Appendix F Implementation Details in the Main Experiments
@@
Here we provide a detailed account of the implementation specifics from the main experiments in §6. For Self-Consistency, we generate 32 samples and choose the major voting answer as the final prediction. For BoN, we set the temperature to 0.7 to control the diversity and choose the best answer from 32 samples. For BoN Weighted, we normalize the RM’s scoring and use this score as a weight to conduct a weighted vote among different answers, selecting the final prediction. For MCTS, we set the rollout number to 16, the width to 5, the max depth to 5, and the explore weight to 0.1. For Beam Search, we set the Beam numbers to 8, the beam width to 5, and the max depth to 5. For our method, in Step 1, we generate 16 samples with a temperature setting of 0.7 and stop the algorithm if the answer count is less than 2. In Step 2, for ORM, we select the top-1 path, for PRM, we select the top-2 paths. In Step 3, we set the width (i.e. the children numbers for each node) to 8 for ORM, 4 for PRM, and the max depth to 3. We release all of the prompts we use in the accompanying software package.
Appendix G Additional Results of the Main Experiments
Results on More Datasets
To test whether our method remains effective on more challenging tasks, we further compared the performance of our method with other baselines on the OlympiadBench dataset (He et al., 2024) in Figure 22. The results demonstrate the effectiveness of OCTS.
Results on More Models
We repeat the main experiments on the Llama3.1-8B model and report the result in Table 7 (here we sample 200 questions from each dataset). Our method also demonstrates the best performance on this model across three different tasks.
Ablation Study
To verify the effectiveness of each step, here we conducted ablation experiments on a combination of Qwen2.5-3B + ORM, using 200 samples each from GSM8k and MATH. The experimental settings are as follows:
  • -Exploration: Disable the early stopping mechanism in Step 1;
@@ Appendix G Additional Results of the Main Experiments
  • -Selection: Eliminate the clustering operation and use the score of each path instead of cluster scores for selection (similar to MCTS);
@@
  • -Expansion: Cancel the operation of directly generating the remaining steps, and instead generate intermediate nodes layer by layer (similar to MCTS and Beam).
@@
Table 8 shows the result of the ablation study. Removing each component leads to a decline in performance. Specifically, although removing exploration causes only a small drop, its inclusion not only improves performance but also reduces inference time.
@@
(a) ORM
(b) PRM
Figure 22: Performance comparison on OlympiadBench (200 samples).
Methods GSM8k MATH OlympiadBench
CoT 84.5 38.0 10.5
Self-Consistency 91.0 56.5 16.0
Best-of-N + ORM 90.5 47.0 17.6
+ PRM 95.0 62.0 23.0
BoN Weighted + ORM 88.5 52.5 19.8
+ PRM 94.0 61.5 24.0
MCTS + ORM 90.0 43.0 12.5
+ PRM 95.0 57.0 19.0
Beam Search 94.0 55.5 14.5
Ours + ORM 88.5 49.0 18.0
+ PRM 95.0 67.0 26.0
Table 7: Performance comparison on Llama3.1-8B, the best results are highlighted in bold.
Methods GSM8k MATH
Ours 91.5 72.5
-Exploration 91.0 72.5
-Selection 89.5 65.5
-Expansion 87.5 64.0
Table 8: Results of the ablation Study.

### v2

@@ 6 Conclusion
In this work, we focus on analyzing key factors that influence the reward model’s performance in reasoning tasks. We find that low question difficulty, large sampling number, and high search diversity can lead to issues in RM-based inference, with in-depth explanations provided. To address these issues, we propose CRISP, a cluster-based, prefix-guided inference algorithm that enhances the robustness and efficiency of the reward model. Experimental results demonstrate that our method is effective in enhancing LLM reasoning capabilities.
Reproducibility Statement
We have taken several steps to improve the reproducibility of our research. We offer a detailed account of the parameter settings and prompts used in the experiments, which are outlined in Appendix L. The full experimental code is also uploaded in the supplementary materials. We commit to making all code open source if the paper is accepted.
Appendix A The Use of Large Language Models
Throughout the preparation of this manuscript, a large language model (LLM) was employed to assist exclusively with language refinement. Specifically, the LLM was used for:
  • Grammar and Syntax Improvements: Correcting errors and optimizing sentence structures.
  • Conciseness and Precision: Providing alternative phrasings for brevity and accuracy.
All research concepts, analyses, and conclusions were developed independently by the authors. The LLM’s contributions were limited to linguistic enhancement and did not influence the study’s conceptual content.
Appendix B Limitations & Future Work
While our work provides a thorough investigation of RM behavior during inference, it does not address potential issues that may arise during the training of models. In future work, we aim to extend our study to the training phase of reward models. Understanding how training dynamics (such as reward signal design and data sampling strategies) impact downstream reasoning performance could offer deeper insights and help improve the overall reliability of LLM.
Appendix C Performance of Selected RMs
To demonstrate that the RM issues identified in our experiments in Section §2 are not due to the selected RM’s inherently low discriminative abilities, here we present the performance of our RM. For the two ORMs (e.g. ArmoRM-Llama3-8B and Skywork-Reward-Llama-3.1-8B), we report their performance on RewardBench (Lambert et al., 2024) compared to other baselines in Table 5. For the two PRMs (e.g. Math-Shepherd-Mistral-7B-PRM and Skywork-o1-Open-PRM-Qwen-2.5-7B), we report their performance on ProcessBench (Lambert et al., 2024) compared to other baselines in Table 6. From them, we can get that the performance of these models on relevant benchmarks is comparable to the advanced LLMs (e.g., GPT -4), hence they are representative.
Appendix D Additional Overall Experiments
@@
In addition, we also use the consistency of the final answer output by the policy model itself as the source of the reward, denoted as ‘Self’. The results are demonstrated in Figure 11. We can conclude that: (1) Even with the MCTS framework, the improvement in model reasoning brought by the RM is still minimal, further validating our conclusions in §2. (2) In Skywork and Skyworko1, the average performance of Reward is the best among all scoring functions. Therefore, in the MCTS-related experiments presented in the main text, we default to using it as the scoring function $f$.
Secondly, we focus on math reasoning in the main text, here we repeat our experiments on other types of reasoning tasks. Specifically, for math reasoning, we select another dataset: AQuA (Ling et al., 2017). For commonsense reasoning, we select WinoGrande (WINO) (Sakaguchi et al., 2020) and CSQA (Talmor et al., 2019); For logical reasoning, we select ProofWriter (Tafjord et al., 2021) and ProntoQA (Saparov and He, 2023) The results are demonstrated in Figure 12, 13, 14, 15 and 16. Lastly, we only use discriminative RM in the main text. All of these results are consistent with the conclusion in the main text.
Appendix E Additional Experiments on Question Difficulty Approximation
@@
We classify the problems according to the difficulty levels as outlined in the main text and calculate the above three metrics across different levels of problem difficulty to compare the degree of correlation. The results are illustrated in Figure 17, 18 and 19. We can observe that, comparatively, the Count function is most directly proportional to difficulty. Therefore, we use this function to estimate difficulty when designing the CRISP method in §4.1.
Appendix F Additional Experiments across Different Difficulty Levels
In the main text, we only analyze the impact of question difficulty on the MATH dataset. To demonstrate the generalizability of our conclusions, we repeat this experiment on GSM8K (Cobbe et al., 2021) and Olympiadbench (He et al., 2024). The former dataset contains 8.5K linguistically diverse elementary school math problems designed to evaluate arithmetic reasoning consistency, while the latter is an Olympiad-level bilingual multimodal scientific benchmark. Compared to MATH, the former is simpler, while the latter is more challenging. The results are illustrated in Table 8, 9 and 10. We can observe that the issues identified in Cl.1 are prevalent across various reasoning datasets.
Appendix G Comparison Between Coverage and Accuracy
The changes in accuracy and coverage are shown in Figure 20,21. The results demonstrate that: Regardless of the inference strategy used, the model’s accuracy does not improve as $n$ increases. The accuracy in plateaus beyond a relatively small number of samples (approximately 30). In contrast, the Oracle setting consistently increases, leading to a persistently widening gap between accuracy and coverage.
Appendix H Case Analysis of Sampling Numbers Experiment
Appendix I Cause Analysis of Temperature-Induced Accuracy Drop
We further conduct statistical analyses to uncover the reasons for this issue. For each $T$, we calculate the information entropy of incorrect answers across 16 samplings and report the distribution over 200 questions in Figure 7, 24. As the temperature rises, the entropy for both models shows a gradually increasing trend, hence, the distribution of these negative samples becomes more random. This indicates that the policy model generates a greater number of low-frequency incorrect answers at higher temperatures. According to Cl.2, RM struggles to differentiate these negative examples from correct ones, leading to lower inference accuracy. This result not only elucidates the reasons behind the subpar performance of BoN under high diversity conditions but also further corroborates the inverse long-tail phenomenon of the RM.
Appendix J Diversity Experiment on Exploration Constant
@@ Appendix I Cause Analysis of Temperature-Induced Accuracy Drop
Appendix K Theoretical Analysis of CRISP Method
In this section, we present a theoretical analysis of the clustering strategy (i.e., State Aggregation module + Reward Evaluation module) within the CRISP method, as it serves as the core component of the entire approach.
Assume we have sampled $n$ paths, where each answer $a_{i}$ corresponds to a reward $r_{i}$, and $f_{i}$ is the frequency of $a_{i}$. In Cl.2, we observe that RM tends to assign a higher $r_{i}$ to an incorrect $a_{i}$ with lower $f_{i}$, sometimes even exceeding the score of the highest-scoring correct example, leading to an incorrect final answer. Our CRISP’s clustering method incorporates frequency $f_{i}$ as a factor into the new reward scores $r^{\prime}_{i}$ to mitigate this issue:
$r^{\prime}_{i}=\sum_{a_{k}=a_{i}}r_{k}=f_{i}\cdot\overline{r_{i}}$ (7)
where $\overline{r_{i}}$ represents the average score of the cluster to which $a_{i}$ belongs. Suppose $a_{j}$ is the top-scored negative answer, we have:
$\frac{r^{\prime}_{i}}{r^{\prime}_{j}}=\frac{f_{i}}{f_{j}}\cdot\frac{\overline{r_{i}}}{\overline{r_{j}}}$ (8)
where $\overline{r_{i}}$ represents the average score of the cluster to which $a_{i}$ belongs. Although $\overline{r_{i}}<\overline{r_{j}}$, as long as $\frac{f_{i}}{f_{j}}>\frac{\overline{r_{j}}}{\overline{r_{i}}}$, we have $r^{\prime}_{j}>r^{\prime}_{i}$. According to Figure 5, when n=128, in most cases, $f_{j}<3$, which is a very small value. Therefore, in most cases, there exists $f_{i}\gg f_{j}$, such that $r^{\prime}_{j}>r^{\prime}_{i}$, reducing the score ranking of these negative examples.
In summary, our CRISP method reduces the tendency of the RM to assign excessively high scores to low-frequency negative examples, thereby increasing the probability of selecting the correct path. It performs better when the generative model samples the correct answer more frequently (i.e., $f_{i}\gg f_{j}$).
Appendix L Implementation Details in the Main Experiments
Here, we provide a detailed account of the implementation specifics from the main experiments:
@@
For Self-Consistency, we generate 32 samples and choose the major voting answer as the final prediction. For BoN, we set the temperature to 0.7 to control the diversity and choose the best answer from 32 samples. For BoN Weighted, we normalize the RM’s scoring and use this score as a weight to conduct a weighted vote among different answers, selecting the final prediction. For MCTS, we set the rollout number to 16, the width to 5, the max depth to 5, and the explore weight to 0.1. For Beam Search, we set the Beam numbers to 8, the beam width to 5, and the max depth to 5.
For our method, we generate 16 samples with a temperature setting of 0.7 in the first iteration. In subsequent iterations, we set the sampling numbers to 8 for ORM, 4 for PRM, and the max depth to 3. In prefix extraction, for ORM, we select the top-1 path, for PRM, we select the top-2 paths. Tables 11, 12, 13 and 14 present the key experimental results demonstrating our exploration of different hyperparameter configurations. We can get the following key takeaways:
  • For the threshold, we should set a larger value for simpler tasks (such as GSM8K) and a smaller value for more difficult tasks (such as MATH). This is because a larger threshold makes our method equivalent to SC in more cases, and as shown in Cl.1, SC performs better than RM-based inference on simpler tasks.
  • For the max steps m and top k, we should set them to higher levels for simpler tasks, while for more difficult tasks, they should be set to moderate values, without being too high (e.g., m = 3 and k = 2). This is because excessively large parameters introduce higher sampling diversity, which, as shown in Cl.3, results in more high-quality negative examples. This can particularly degrade performance on more difficult tasks.
  • For the sampling numbers, we find that increasing n does not continuously lead to better performance. Therefore, in the main paper, we set $N$ to a moderate value of 16 to control the cost.
For the evaluation data, we sample 500 questions from GSM8K and MATH-500, while sampling 200 questions from OlympiadBench. We release the prompts we use in Table 19, 20, 21, 22, 23 and 24. All experiments were conducted on NVIDIA A100 GPUs.
Appendix M Ablation Study
To verify the effectiveness of each module of CRSIP, we conduct ablation experiments on different modules in it. The experimental settings are as follows:
  • w/o Termination: Disable the early termination condition based on the number of clusters;
@@
  • w/o Aggregation: Eliminate the clustering operation and use the score of each path instead of cluster scores for selection (similar to MCTS);
@@
  • w/o Prefixing: Cancel the operation of directly generating the remaining steps according to the prefix set, and instead generate intermediate nodes layer by layer (similar to MCTS and Beam).
@@
Figure 28 and Table 18 show the result of the ablation study. Removing each component leads to a decline in performance. Specifically, although w/o termination causes only a small drop, its inclusion not only improves performance but also reduces inference time.
Appendix N Additional Experiments on Dataset Difficulty Splits
We introduce the difficulty level from the original MATH-500 dataset (Hendrycks et al., 2021), which is independent of any specific model, in order to more objectively compare the performance of different paradigms across varying difficulty levels. The results are shown in Table 7. The results show that the question difficulty in our findings is actually independent of the specific model.
Appendix O Further Discussion on the Main Experiments
We provide the significance test results for the main experiments to demonstrate that our method consistently improves performance. Specifically, we repeat the experiments on the MATH dataset using Qwen-2.5-3B + Skyworko1 for five runs. The results are in Table 15. Based on the results, we conduct t-test experiments (see Table 16) and calculate confidence intervals (see Table 17). The results demonstrate that our method consistently and significantly outperforms the other baselines.
@@
Table 7: Comparison of performance across difficulty levels split by the MATH dataset (Qwen2.5-3B).
Method Level 1 Level 2 Level 3 Level 4 Level 5
SC 0.93 0.80 0.76 0.61 0.37
BoN (ORM) 0.86 0.72 0.73 0.62 0.38
BoN (PRM) 0.79 0.71 0.71 0.60 0.46
@@
Figure 17: The correlation between output length and the question difficulty.
Figure 18: The correlation between the count of answers and the question difficulty.
Figure 19: The correlation between the count of no answers and the question difficulty.
Figure 20: BoN performance across different sampling numbers.
Figure 21: MCTS performance across different sampling numbers.
(a) BoN
(b) MCTS
Figure 23: Frequency statistics of the highest-scored negative responses in MCTS.
@@
Figure 26: Performance of BoN inference across different sampling temperatures (Llama3.1-8B).
(a) Tree width
(b) Tree depth
Figure 27: MCTS inference performance under different tree structures (PRM).
Algorithm 1 Clustered Reward Integration with Stepwise Prefixing
1:Policy model $\mathcal{M}$, reward score $f$, question $q$, max steps $m$, sampling numbers $n$, top-$k$ parameter $k$
[…]
7:while $i<n$ do
8:  if $i=0$ then
9:   $\mathcal{R}\leftarrow\mathcal{M}(q,n)$ $\triangleright$ Generate $n$ initial responses
10:   if $|\operatorname{Cluster}(\mathcal{R})|=1$ then
11:     return $\mathcal{R}[0]$ $\triangleright$ Early exit if only one cluster
12:   end if
13:  else
14:   $\mathcal{R}_{\text{top}}\leftarrow\left\{\arg\max_{r\in\mathcal{C}_{j}}f(r)\,\middle|\,\mathcal{C}_{j}\in\mathcal{C}_{\text{top}}\right\}$
15:   $\mathcal{P}\leftarrow\{r[{:}i{+}1]\mid r\in\mathcal{R}_{\text{top}}\}$ $\triangleright$ Truncate top responses
16:   $\mathcal{R}\leftarrow\mathcal{R}\cup\mathcal{M}(q,n,\mathcal{P})$ $\triangleright$ Decode more based on prefixes
17:  end if
18:  $\mathcal{C}\leftarrow\operatorname{Cluster}(\mathcal{R})$ $\triangleright$ Cluster current responses
19:  for all $\mathcal{C}_{j}\in\mathcal{C}$ do
20:   $\mathcal{F}(\mathcal{C}_{j})\leftarrow\sum_{x\in\mathcal{C}_{j}}f(x)$ $\triangleright$ Assign cluster-wise reward
21:  end for
[…]
23:  $i\leftarrow i+1$
24:end while
25:return $\mathcal{R}_{\text{top}}[0]$
Table 11: Performance comparison under different sampling numbers NN (Qwen2.5-3B + Skywork + MATH).
Methods N = 16 N = 32 N = 64 N = 128
SC 0.65 0.64 0.64 0.65
BoN 0.64 0.67 0.66 0.63
MCTS 0.64 0.65 0.67 0.65
Ours 0.69 0.72 0.69 0.69
Table 12: Performance comparison under different top-k values (Qwen2.5-3B + Skywork).
Datasets k = 1 k = 2 k = 4
GSM8K 0.90 0.91 0.91
MATH 0.73 0.74 0.70
Table 13: Performance comparison under different cluster threshold (Qwen2.5-3B + Skywork).
Datasets Threshold = 2 Threshold = 3 Threshold = 4 Threshold = 5
GSM8K 0.90 0.88 0.91 0.92
MATH 0.73 0.72 0.69 0.70
Table 14: Performance comparison under different max steps $m$ (Qwen2.5-3B + Skywork).
Datasets m = 2 m = 3 m = 5
GSM8K 0.90 0.90 0.92
MATH 0.70 0.73 0.69
Figure 28: Results of our ablation study on different datasets.
Table 15: Five-run results of our main experiments on MATH (Qwen2.5-3B + Skyworko1).
Methods Round 1 Round 2 Round 3 Round 4 Round 5
BoN 0.61 0.64 0.63 0.70 0.68
Weighted SC 0.63 0.71 0.67 0.71 0.71
MCTS 0.71 0.72 0.73 0.73 0.68
Beam Search 0.72 0.70 0.72 0.70 0.65
Ours 0.76 0.78 0.77 0.78 0.75
Table 16: $t$-test results for the main experiments on MATH (Qwen2.5-3B + Skyworko1).
Comparison t-statistic p-value
BoN vs Ours -6.859220 0.002365
Weighted SC vs Ours -5.360510 0.005844
MCTS vs Ours -10.590300 0.000450
Beam Search vs Ours -6.390100 0.003079
Table 17: Confidence intervals, results for the main experiments on MATH (Qwen2.5-3B + Skyworko1).
Method Mean Variance 95% Confidence Interval
BoN 0.652 0.0011 0.652 ± 0.029
Weighted SC 0.686 0.0012 0.686 ± 0.030
MCTS 0.714 0.0005 0.714 ± 0.020
Beam Search 0.698 0.0007 0.698 ± 0.023
Ours 0.768 0.0002 0.768 ± 0.012
Table 18: Results of our ablation study on different reward models.
Method ORM PRM
Ours 0.73 0.78
-w/o Termination 0.72 0.76
-w/o Aggregation 0.71 0.75
-w/o Prefixing 0.64 0.72
(a) Time Consumption Comparison (s)
(b) Token Consumption Comparison
Figure 29: Results of our cost analysis.
Figure 30: Compute-return curve on GSM8K (Qwen2.5-3B + Skywork).
Table 19: Prompts used to sample reasoning paths on the GSM8K dataset.
Prompt
[…]
# Question:
Mr. Ruther sold $\frac{3}{5}$ of his land and had 12.8 hectares left. How much land did he have at first?
# Reasoning:
Step 1: Mr. Ruther is left with $1-\frac{3}{5}=\frac{2}{5}$ of his land.
Step 2: Since $\frac{2}{5}$ equals 12.8 hectares, then $\frac{1}{5}=\frac{12.8}{2}=6.4$ hectares.
Step 3: Total land = $6.4\times 5=32$ hectares.
Step 4: The answer is: $\boxed{32}$
# Question:
The Doubtfire sisters are driving home with 7 kittens adopted from the local animal shelter when their mother calls to inform them that their two house cats have just had kittens. She says that Patchy, the first cat, has had thrice the number of adopted kittens, while Trixie, the other cat, has had 12. How many kittens does the Doubtfire family now have?
# Reasoning:
Step 1: Patchy has had $3\times 7=21$ kittens.
Step 2: Trixie has had 12 kittens. Total from both cats = $21+12=33$.
Step 3: Total kittens including adopted = $7+33=40$.
Step 4: The answer is: $\boxed{40}$
# Question:
After transferring to a new school, Amy made 20 more friends than Lily. If Lily made 50 friends, how many friends do Lily and Amy have together?
# Reasoning:
Step 1: Amy made $50+20=70$ friends.
Step 2: Total friends = $70+50=120$.
Step 3: The answer is: $\boxed{120}$
# Question:
{current question}
# Reasoning:
Table 20: Prompts used to sample reasoning paths on the MATH dataset.
Prompt
Please act as a math teacher and give step-by-step solutions to the user’s questions. At the final step, a conclusive answer is given in the format of ”The answer is: <ANSWER>.”, where <ANSWER> should be a numeric answer.
# Question:
How many 3-letter words can we make from the letters A, B, C, and D, if we are allowed to repeat letters, and we must use the letter A at least once? (Here, a word is an arbitrary sequence of letters.)
# Reasoning:
Step 1: There are $4^{3}$ three-letter words from A, B, C, and D, and there are $3^{3}$ three-letter words from just B, C, and D.
Step 2: There must, then, be $4^{3}-3^{3}=64-27=\boxed{37}$ words from A, B, C, and D containing at least one A.
Step 3: The answer is: $\boxed{37}$
# Question:
In the diagram, square $ABCD$ has sides of length 4, and $\triangle ABE$ is equilateral. Line segments $BE$ and $AC$ intersect at $P$. Point $Q$ is on $BC$ so that $PQ$ is perpendicular to $BC$ and $PQ=x$.
# Reasoning:
Step 1: Since $\triangle ABE$ is equilateral, we know that $\angle ABE=60^{\circ}$.
[…]
# Question:
Find the positive real number(s) $x$ such that $\frac{1}{2}(3x^{2}-1)=(x^{2}-50x-10)(x^{2}+25x+5).$ # Reasoning:
Step 1: Write $a=x^{2}-50x-10$ and $b=x^{2}+25x+5$.
Step 2: Then the equation given becomes $\frac{a+2b-1}{2}=ab,$ so 0=2aba2b+1=(a1)(2b1)0=2ab-a-2b+1=(a-1)(2b-1).
Step 3: Then $a-1=x^{2}-50x-11=0$ or $2b-1=2x^{2}+50x+9=0$.
Step 4: The former has a positive root, $x=\boxed{25+2\sqrt{159}}$, while the latter does not.
Step 5: The answer is: $\boxed{25+2\sqrt{159}}$
# Question:
{current question}
# Reasoning:
Table 21: Prompts used to sample reasoning paths on the Olympiadbench dataset.
Prompt
[…]
# Question:
Let $T$ be a rational number. Compute $\sin^{2}\frac{T\pi}{2}+\sin^{2}\frac{(5-T)\pi}{2}$.
# Reasoning:
Step 1: Note that $\sin\frac{(5-T)\pi}{2}=\cos\left(\frac{\pi}{2}-\frac{(5-T)\pi}{2}\right)=\cos\left(\frac{T\pi}{2}-2\pi\right)=\cos\frac{T\pi}{2}$.
Step 2: Thus the desired quantity is $\sin^{2}\frac{T\pi}{2}+\cos^{2}\frac{T\pi}{2}=\boxed{1}$.
Step 3: The answer is: $\boxed{1}$
# Question:
Let $T=11$. Compute the value of $x$ that satisfies $\sqrt{20+\sqrt{T+x}}=5$.
# Reasoning:
Step 1: Squaring both sides gives $20+\sqrt{T+x}=25$, so $\sqrt{T+x}=5$.
Step 2: Squaring again gives $T+x=25$, so x=25T=14x=25-T=14.
Step 3: The answer is: $\boxed{14}$
# Question:
The sum of the interior angles of an $n$-gon equals the sum of the interior angles of a pentagon plus the sum of the interior angles of an octagon. Compute $n$.
# Reasoning:
Step 1: The sum of interior angles of an $n$-gon is $180^{\circ}(n-2)$.
Step 2: A pentagon has sum $180^{\circ}(5-2)=540^{\circ}$, and an octagon has sum $180^{\circ}(8-2)=1080^{\circ}$.
Step 3: So 180(n2)=540+1080=1620180(n-2)=540+1080=1620, hence n2=9n-2=9, so $n=11$.
Step 4: The answer is: $\boxed{11}$
# Question:
{current question}
# Reasoning:
Table 22: Prompts used to sample reasoning paths on the CSQA dataset.
Prompt
Please act as a commonsense teacher and solve the commonsense reasoning problem step by step.
# Question:
Google Maps and other highway and street GPS services have replaced what?
# Options:
(A) atlas  (B) mexico  (C) countryside  (D) united states  (E) oceans
# Reasoning:
Step 1: Electronic maps and GPS services are the modern version of paper atlas.
Step 2: In that case, the atlas have been replaced by Google Maps and other highway and street GPS services.
Step 3: The answer is: A
# Question:
You can share files with someone if you have a connection to a what?
# Options:
(A) freeway  (B) radio  (C) wires  (D) computer network  (E) electrical circuit
# Reasoning:
Step 1: Files usually can be stored in the computers.
Step 2: In that case, we can share them over the Internet.
Step 3: Thus, if we connect to a computer network, we can share the file with others.
Step 4: The answer is: D
# Question:
The fox walked from the city into the forest, what was it looking for?
# Options:
(A) pretty flowers  (B) hen house  (C) natural habitat  (D) storybook  (E) dense forest
# Reasoning:
Step 1: Since the fox walk from the city into the forest, he may looks for something in the forest but not in the city.
Step 2: From all of the options, the natural habitat are usually away from cities.
Step 3: The answer is: C
# Question:
{current question}
# Options:
{current options}
# Reasoning:
Table 23: Prompts used to sample reasoning paths on the SIQA dataset.
Prompt
Please act as a commonsense teacher and solve the commonsense reasoning problem step by step.
# Question:
Quinn wanted to help me clean my room up because it was so messy. What will Quinn want to do next?
# Options:
(A) Eat messy snacks  (B) help out a friend  (C) Pick up the dirty clothes
# Reasoning:
Step 1: Quinn wants to clean the room up.
Step 2: Picking up the dirty clothes is one way to clean the room.
Step 3: Thus, Quinn will want to pick up the dirty clothes next.
Step 4: The answer is: C
# Question:
Sydney had so much pent up emotion, they burst into tears at work. How would Sydney feel afterwards?
# Options:
(A) affected  (B) like they released their tension  (C) worse
# Reasoning:
Step 1: Crying is often a way to release tension.
Step 2: Sydney burst into tears at work.
Step 3: Thus, she would release the tension.
Step 4: The answer is: B
# Question:
Their cat kept trying to escape out of the window, so Jan placed an obstacle in the way. How would Jan feel afterwards?
# Options:
(A) scared of losing the cat  (B) normal  (C) relieved for fixing the problem
# Reasoning:
Step 1: The cat tried to escape so Jan needed to stop it to avoid losing the cat.
Step 2: Jan placed an obstacle in the way so the cat could not escape.
Step 3: The problem has been solved.
Step 4: Thus, Jan will feel relieved for fixing the problem.
Step 5: The answer is: C
# Question:
{current question}
# Options:
{current options}
# Reasoning:
Table 24: Prompts used to sample reasoning paths on the LogiQA dataset.
Prompt
Please act as a logical teacher and reason step by step to solve the logical reasoning problem.
# Context:
There are 90 patients with a disease T that is very difficult to treat and has taken the same routine drug. The patients were divided into two equal groups. The first group was given an experimental drug W, which is used to treat T, and the second group was given a placebo without W. Statistics ten years later showed that 44 people died in both groups, so the experimental drug was ineffective.
# Question:
Based on the above information, which of the following options, if correct, will best weaken the above argument?
# Options:
(A) Among the patients who died above, the average year of death in the second group was two years earlier than that in the first group.  (B) Among the patients who died, the average life span of the second group was two years younger than that of the first group.  (C) Among the above-mentioned living patients, the condition of the second group was more serious than that of the first group.  (D) Among the above-mentioned living patients, those in the second group were older than those in the first group.
# Reasoning:
Step 1: Analyzing each option: A suggests drug W might extend life since the average death year in the drug W group is later than the placebo, directly challenging the drug’s perceived ineffectiveness.
Step 2: B, similar to A, implies longer life in the drug W group but doesn’t directly link to post-treatment lifespan.
Step 3: C indicates drug W may reduce disease severity but doesn’t address lifespan or mortality, the main focus.
Step 4: D, about age differences, lacks direct relevance to drug effectiveness.
Step 5: Therefore, A most effectively weakens the argument against drug W’s effectiveness.
Step 6: The answer is: A
# Question:
{current question}
# Options:
{current options}
# Reasoning: